# Beyond the Turn-Based Game: Enabling Real-Time Conversations with Duplex Models

### v1

@@ Abstract
Xinrong Zhang, Yingfa Chen, Shengding Hu, Xu Han, Zihang Xu, Yuanwei Xu, Weilin Zhao Zhiyuan Liuthanks: Corresponding author: Xu Han and Zhiyuan Liu, Maosong Sun NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China zxr19@mails.tsinghua.edu.cn, {hanxu2022,liuzy}@tsinghua.edu.cn
@@ 1 Introduction
Despite ongoing advancements, interactions with LLMs often fail to provide users with human-like interaction experience Hill et al. (2015); Mou and Xu (2017); Zhou et al. (2023). One reason is the turn-based nature of current chatbot implementations (Skantze, 2021), which is different from human conversations where there are many overlaps, interruptions, and silences (Zimmerman and West, 1996). Current human-LLM interactions necessitate that one participant remains entirely idle while the other generates responses. Interruptions are manually triggered with a “stop” button or by saying certain keywords, resulting in conspicuously artificial communication. In human conversations, participants simultaneously process incoming information and formulate responses, often in overlapping and interleaved contexts, thus allowing each other to interrupt or be interrupted.
To address this limitation, we introduce the concept of duplex models. Duplex models emulate human cognitive processes by synthesizing responses internally while simultaneously attending to incoming user inputs, akin to a person thinking while listening as well as speaking while observing. However, present autoregressive models face substantial challenges in adopting a duplex configuration, as they must process and encode a complete input message before generating any tokens, resulting in a turn-based conversation. Considering this, we propose a framework for quickly converting current LLMs into duplex models by processing queries and responses pseudo-simultaneously without significant alternations to their architectures.
@@ 3 Supervised Fine-Tuning Duplex Dataset
Example Type # Dialogues Avg. # Slice Pairs Avg. # Tokens
@@
Table 1: The statistics of Duplex-UltraChat. The tokens are produced by the tokenizer of our MiniCPM-duplex.
@@ 3.2 Dialogues with Interruptions
Forced interruptions are when users directly speak out their next sentence regardless of the status of the assistant. To generate such data, we randomly choose a location in an assistant message, discard the remaining part of the message, and insert a new user input at that location. We prefix the user input with one of the 11 pre-defined transitional sentences (see Appendix A.1). This input is rewritten by ChatGPT to ensure a natural and varied transition. The target output is idle tokens because the assistant is expected to terminate its current response.
@@ 4.1 Training
The training of MiniCPM-duplex uses the following hyperparameters: $10^{-3}$ maximum learning rate, warmupstableexp (Hu et al., 2024) learning rate scheduler, a batch size of 800, and a maximum length of 4,096. We train for 10,000 steps on 40 NVIDIA A100 GPUs for 36 hours.
@@ 4.3.1 GPT-4 Evaluation
To evaluate the multi-turn dialogue ability of MiniCPM-duplex, we benchmark it on MT-Bench (Zheng et al., 2024) with GPT-4 as the judge.
To mimic real-time scenarios, we chunk each instruction in MT-Bench into multiple 4-6 word slices and feed one slice at a time. Then we concatenate all output segments from the duplex model to form the final output. For the traditional model, we directly feed the entire prompt to the model.
@@ 5 Results
Table 4 shows the GPT-4 evaluation results on MT-Bench. MiniCPM-duplex is slightly inferior to MiniCPM mainly due to that MiniCPM-duplex tends to generate shorter responses. GPT-4 favors longer responses, whereas users prefer chat models that give concise answers.
@@ 6.1 Analysis
The superior performance of the duplex model is mainly due to its underlying receive/generate mechanism. Rather than strictly turn-based dialogue where users must explicitly signal the beginning and end of messages, duplex models behave more like human beings. Besides, the duplex model has learned when to speak at the fine-tuning stage on the Duplex-UltraChat, which makes it more human-like. Such ability is essential in passing a non-turn-based version of the Turing test, which is a more realistic test for whether a machine can be indistinguishable from humans.
@@ 7.2 Dialogue Models
Chat-based models have gained widespread popularity since the release of ChatGPT. Some notable chat-based LLMs include the Claude series (Anthropic, 2023, 2024), Qwen series (Qwen, 2024), the Mistral series (Jiang et al., 2023) and and LLaMa series (Touvron et al., 2023), among others. Most of these models, especially open-sourced ones, are purely text-based.

### v2

@@ Abstract
Xinrong Zhang1, Yingfa Chen1, Shengding Hu1, Xu Han1, Zihang Xu1, Yuanwei Xu3 Weilin Zhao1, Zhiyuan Liu2,1thanks: Corresponding author: Xu Han and Zhiyuan Liu, Maosong Sun1 1NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China. 2Quan Cheng Laboratory, Jinan, China. 3Modelbest Inc. zxr19@tsinghua.org.cn, {hanxu2022,liuzy}@tsinghua.edu.cn
@@ 1 Introduction
Despite ongoing advancements, interactions with LLMs often fail to provide users with human-like interaction experience Hill et al. (2015); Mou and Xu (2017); Zhou et al. (2023). One reason is the turn-based nature of current chatbot implementations (Skantze, 2021), which is different from human conversations where there are many overlaps, interruptions, and silences (Zimmerman and West, 1996).
Current human-LLM interactions require one participant to remain idle while the other generates responses. Interruptions are manually triggered with a “stop” button or certain keywords, resulting in conspicuously artificial communication. In human conversations, participants simultaneously process incoming information and formulate responses, often in overlapping and interleaved contexts, thus allowing each other to interrupt or be interrupted.
We introduce the concept of duplex models to address this limitation. Duplex models emulate human cognitive processes by synthesizing responses internally while simultaneously attending to incoming user inputs, akin to a person thinking while listening as well as speaking while observing. However, present autoregressive models face substantial challenges in adopting a duplex configuration, as they must process and encode a complete input message before generating any tokens, resulting in a turn-based conversation.
Considering this, we propose a framework for quickly converting current LLMs into duplex models by processing queries and responses pseudo-simultaneously without significant alternations to their architectures.
@@ 3 Supervised Fine-Tuning Duplex Dataset
Example Type Interruption # Dialogues Avg. # Slice Pairs Avg. # Tokens
@@
Table 1: The statistics of Duplex-UltraChat. The tokenizer of our MiniCPM-duplex produces the tokens.
@@ 3.2 Dialogues with Interruptions
Forced interruptions are when users directly speak out their next sentence regardless of the status of the assistant. To generate such data, we randomly choose a location in an assistant message, discard the remaining part of the message, and insert a new user input at that location. We prefix the user input with one of the 11 pre-defined transitional sentences (see Appendix A.1). ChatGPT rewrites this input to ensure a natural and varied transition. The target output is idle tokens because the assistant is expected to terminate its current response.
@@ 4.1 Training
The training of MiniCPM-duplex uses the following hyperparameters: $10^{-3}$ maximum learning rate, Warmup-Stable-Decay (Hu et al., 2024) learning rate scheduler, a batch size of 800, and a maximum length of 4,096. We train for 10,000 steps on 40 NVIDIA A100 GPUs for 36 hours.
@@ 4.3.1 GPT-4 Evaluation
To evaluate the multi-turn dialogue ability of MiniCPM-duplex, we benchmark it on MT-Bench (Zheng et al., 2024) and MT-Eval(kwan2024mt) with GPT-4 as the judge.
To mimic real-time scenarios, we chunk each instruction in MT-Bench and MT-Eval into multiple 4-6 word slices and feed one slice at a time. Then we concatenate all output segments from the duplex model to form the final output. For the traditional model, we directly feed the entire prompt to the model.
@@ 4.4 Interactive Demo
Score MiniCPM MiniCPM-duplex
Refinement-multi 5.87 5.78
Expansion-multi 6.31 5.80
Follow-up-multi 8.48 8.56
Avg. 6.88 6.71
Table 5: MT-Eval results of MiniCPM and MiniCPM-duplex. Higher is better.
@@ 5 Results
Table 4 and Table 5 show the GPT-4 evaluation results on MT-Bench and MT-Eval, respectively. MiniCPM-duplex is slightly inferior to MiniCPM mainly due to that MiniCPM-duplex tends to generate shorter responses. GPT-4 favors longer responses, whereas users prefer chat models that give concise answers.
@@ 6.1 Analysis
The superior performance of the duplex model is mainly due to its underlying receive/generate mechanism. Rather than strictly turn-based dialogue where users must explicitly signal the beginning and end of messages, duplex models behave more like human beings. Besides, the duplex model has learned when to speak at the fine-tuning stage on the Duplex-UltraChat, which makes it more human-like. Such ability is essential in passing a non-turn-based version of the Turing test, which is a more realistic test for whether a machine can be indistinguishable from humans (barnaud2017perceptuo).
@@ 7.2 Dialogue Models
Chat-based models have gained widespread popularity since the release of ChatGPT. Some notable chat-based LLMs include the Claude series (Anthropic, 2023, 2024), Qwen series (Qwen, 2024), the Mistral series (Jiang et al., 2023) and LLaMa series (Touvron et al., 2023), among others. Most of these models, especially open-sourced ones, are purely text-based.
@@ 8 Conclusion
Acknowledge
This work was supported by the National Key R&D Program of China (No.2022ZD0116312), Quan Cheng Laboratory (Grant No.QCLZD202301), and the National Natural Science Foundation of China (No.623B2065).